# Laura Barros        Aiko Schmeißer

### v1

@@ 2.2 Local councilors
Brazil has a highly decentralized political system, with municipal governments managing substantial budgets to provide public services, such as local infrastructure, transportation, education, and health care. Each of Brazil’s 5,568 municipalities is governed by a mayor (prefeito) and a council of local legislators (vereadores). [^6] This number excludes the insular district of Fernando de Noronha and the Federal District which do not take part in local elections. Every four years, during the municipal elections, citizens cast one vote for a mayoral candidate and one vote for a council candidate. [^7] Voting for a party instead of an individual candidate for councilor is also allowed. Local elections are held at the same time for all municipalities across the country, usually in October. Elections for the President, governors, and federal and state legislators also take place every four years but are staggered by two years relative to local elections. Local councilors are elected based on a system of open-list proportional representation and serve four-year terms, with no term limits. The size of local councils varies depending on population, with a minimum of 9 councilors elected in municipalities with up to 15,000 inhabitants and a maximum of 55 councilors elected in municipalities with more than 8 million inhabitants. Local legislators are responsible for drafting laws, voting bills, proposing public works, and monitoring the actions of the executive. In addition, they must approve the municipal budgets and thereby have great influence over the delivery of public goods.
Working as a local councilor can be financially rewarding. The wages of councilors are substantial on average – amounting to about 2.5 times the average wage in their municipalities. Councilor wages also vary strongly by municipality size – being capped at 20% of state deputies’ wages in municipalities with less than 10,000 inhabitants and at 75% in municipalities with more than 500,000 inhabitants. Importantly, serving as a councilor is not a full-time activity and allows legislators to retain their private-sector jobs after being elected. On average, legislators need to be present in the council only four days per month (Ferraz and Finan, 2011). Besides the salaries received by elected councilors, patronage channels can also generate returns for candidates who are not elected. Colonnelli et al. (2020) find that non-elected councilor candidates are about 15 percentage points more likely to be employed in the public sector and receive 40% higher earnings in the formal economy if their supported party comes into power.
@@
How large are the financial costs associated with candidacies for local council? To finance their campaigns, candidates rely on party funds, financial contributions from individuals and firms, as well as own resources. [^10] Campaign spending limits were introduced in 2015 (Avis et al., 2022). Besides banning financial contributions from private firms, the electoral law established spending limits for political campaigns depending on municipalities’ population. Importantly, since 2019, candidates for local council are restricted to finance only up to 10% of the spending limit for their municipality with their own resources to ensure more equitable running conditions. We collected data from the Superior Electoral Court (Tribunal Superior Eleitoral, TSE) on all campaign contributions received by councilor candidates in the electoral cycles 2004 to 2020. Appendix Table A.1 presents detailed summary statistics for each year. Across all years, the median candidate raised a total of 1,587 BRL (392 USD), with only 168 BRL (34 USD) coming from own resources. Only about half of all candidates use any private funds for the campaign. Thus, the majority of resources financing candidates’ campaigns do not originate from their own money, suggesting that the financial barriers to entering politics as a councilor candidate are limited.
@@ 2.3 Theoretical considerations
We now introduce a simple economic framework to illustrate how a layoff affects an individual’s likelihood of becoming a party member or councilor candidate. [^11] For economic models of candidate selection, see, for example, Black (1972), Caselli and Morelli (2004), and Messner and Polborn (2004). Consider an individual who decides how to divide her non-leisure time between pursuing a career in the private sector or in politics. Let $h^{P}$ denote the share of time devoted to politics. If the individual does not engage in politics ($h^{P}=0$), she earns a market hourly wage of $w^{M}$. When entering politics ($0<h^{P}\leq 1$), she foregoes a share of her earnings that is proportional to $h^{P}$. [^12] Since neither party membership nor local councilor is a full-time political commitment, we consider the possibility of simultaneously working in the private sector, albeit with a reduced number of hours. At the same time, the political engagement provides her with an array of expected benefits $b^{P}$. These include monetary rewards, such as councilor salaries when elected to office (Ferraz and Finan, 2011) or public sector employment opportunities resulting from political connections to the winning party or candidate (Brollo et al., 2017; Colonnelli et al., 2020; Barbosa and Ferreira, 2023), as well as non-monetary rewards, such as ego rents or prosocial benefits (Gulzar and Khan, 2024). To engage in politics, the individual incurs entry costs $c^{P}$ (e.g., campaign costs or membership registration fees).
@@
An individual who is currently employed in the private sector decides to enter politics if the benefits of engaging in politics are greater than the benefits of pure market work, i.e., $w^{M}(1-h^{P})+b^{P}h^{P}-c^{P}\geq w^{M}$, or equivalently, if the net benefits are positive:
$\pi(\text{job}=1)=h^{P}(b^{P}-w^{M})-c^{P}\geq 0.$
If the individual is laid off, she loses her earnings opportunities from market work ($w^{M}=0$) and spends no more time in the private sector ($h^{P}=1$). By reducing the private-sector outside options and increasing the time available for politics, the layoff decreases the opportunity costs of entering politics. The decision to engage with politics now depends only on the expected benefits and entry costs:
$\pi(\text{job}=0)=b^{P}-c^{P}\geq 0.$
@@
To determine the effect of a layoff on the likelihood of entering politics, we can compare the net benefits before and after the layoff:
$\pi(\text{job}=0)-\pi(\text{job}=1)=w^{M}h^{P}+b^{P}(1-h^{P}).$
@@
Since $w^{M}\geq 0$, $b^{P}\geq 0$, and $0\leq h^{P}\leq 1$, this difference is weakly positive, suggesting that layoffs induce more individuals to enter politics. The change in net benefits resulting from the layoff has two components: First, it is proportional to the foregone earnings from not spending time $h^{P}$ in the private sector. To test for the relevance of this mechanism, we will analyze effect heterogeneity along individuals’ predicted income loss from the layoff. Second, the effects are predicted to increase in the foregone benefits from not spending time $1-h^{P}$ in politics. To shed light on one of the monetary returns from engaging with politics, we will investigate whether the effects are larger in localities where councilors earn higher salaries. [^13] Besides monetary returns, political involvement may also provide policy payoffs. In the original citizen-candidate models (Osborne and Slivinski, 1996; Besley and Coate, 1997), political candidates represent particular policies and care about winning per se. As Hall (2019) and Thomsen (2017) argue, the policy returns may be higher for ideologically extreme candidates. Thus, we will examine potential heterogeneity in the effects on memberships and candidacies in parties of different ideological positions.
@@
The framework also helps us to think about the implications of layoffs for the quality of those who select into politics. Given that the participation effects of layoffs are predicted to be larger for individuals with higher market wages $w^{M}$ and higher expected benefits $b^{P}$, the consequences for selection depend on how $w^{M}$ and $b^{P}$ are related to an individual’s competence. High-ability individuals tend to receive higher market wages. Thus, a layoff can be seen as a larger shock to their private sector opportunity costs, which leads to a greater increase in their likelihood of entering politics. In addition, the expected benefits may vary with individuals’ competence. For example, the returns of running for office are increasing in the expected probability of winning which may be higher for more competent individuals. Moreover, the expected likelihood of gaining access to a public sector job through political connections may also differ across individuals’ ability levels, either negatively if political support acts as a substitute for individual competence (Colonnelli et al., 2020) or positively if politicians use their discretionary power to appoint higher quality individuals (Brollo et al., 2017).
@@ 3 Data
In our analysis, we combine detailed information on the universe of party members and local political candidates with administrative data on the population of formal employees in Brazil. This section describes the data sources as well as the steps we follow to construct and merge our estimation sample. By having identified records on individuals’ employment histories as well as political outcomes, we can uniquely identify whether individuals’ political behavior changes after they lose their jobs. Additionally, detailed information on individual, party, and municipality characteristics allows us to disentangle potential mechanisms.
3.1 TSE
@@
3.2 RAIS
We obtain matched employer-employee data covering the universe of formal employees between 2004 and 2018 from the Relação Anual de Informações Sociais (RAIS), an administrative database collected annually by the Ministry of Labor and Employment. [^14] In 2015, formally employed individuals accounted for approximately 55% of the Brazilian labor force (Derenoncourt et al., 2021). The database was created in 1975 for statistical and administrative purposes and contains detailed information such as hiring and separation date, average monthly earnings, contracted hours, occupation, firm’s location and industry, and worker’s socio-demographic characteristics. For our analysis, we make use of the identified dataset, containing individuals’ names, taxpayer registry numbers (CPF), as well as firms’ tax identification numbers (Cadastro Nacional da Pessoa Jurídica, CNPJ). This enables us to identify individuals’ complete employment histories over time and across firms.
@@ 3.3 Merging political and labor market data
Appendix Table A.2 compares the characteristics of laid-off individuals in our estimation sample who have a unique and non-unique name, respectively. Those with a unique name are slightly positively selected (9% higher earnings, 8% more years of education, 0.9 percentage points more likely to be a manager), and are more likely to be female (8.1 percentage points) and white (5.3 percentage points). Given that we merge our second political outcome - local candidacies – using the CPF, we have this information for individuals with and without a unique name. Before the layoff, we do not observe differences in the likelihood of running for councilor across both groups. While we focus on the sample of individuals with unique names (in which we have merged political members and candidates) in our baseline specification, we also analyze the external validity of the results by estimating layoff effects on candidacies in the sample of individuals with non-unique names.
4 Empirical Strategy
@@
The main difficulty in estimating the effect of layoffs on individuals’ entry into politics is the fact that layoffs are not random. On the one hand, individual shocks could simultaneously influence both political participation and the likelihood of losing a job, leading to a spurious relationship. On the other hand, individuals’ political behavior could also affect their probability of being dismissed. Colonnelli et al. (2022) show, for instance, that political discrimination affects labor market outcomes in Brazil. To address these identification challenges, we exploit the timing of firm-level mass layoffs as a source of plausibly exogenous variation in individual-level job loss using a difference-in-difference design in a matched sample of workers. The timing of mass layoffs is arguably orthogonal to individual workers’ political behavior and other confounding worker-level shocks, and has been widely used to estimate the effects of job loss on various outcomes. In particular, our empirical strategy closely follows recent papers that study the effect of dismissals in Brazil (Bhalotra et al., 2021; Britto et al., 2022b, a; Amorim et al., 2023). This section starts by describing the steps we follow to construct our estimation sample and then proceeds to present our estimation model.
4.1 Sample selection
@@
We restrict our sample to workers aged 25 to 50 with a full-time (i.e., at least 30 hours per week), open-ended, private-sector job of at least six months tenure. Among those, the treatment group comprises all workers displaced without just cause from a mass layoff firm between 2009 and 2011.161616We exclude individuals who are recalled by their initial employer within five years after the layoff. In cases where individuals experienced multiple mass layoffs between 2009 and 2011, we keep only the first one. This treatment period allows us to study long-run effects over three election cycles (12 years) after the layoff, and also to check for pre-trends over two election cycles (8 years) before the layoff. In our baseline specification, we define a mass layoff firm as a firm with at least 30 workers dismissing at least 30% of its workforce in a given year, and later we show that our results are robust to various thresholds of firm size and layoff share. [^17] We always drop firms from the mass layoff sample if they reallocate under a new identifier, identified as those firms in which at least 30% of their dismissed workers transition to the same new firm identifier in the following year.
@@
Table 1 presents summary statistics for treated and control individuals and shows that both groups are similar in socio-demographic, job, and firm characteristics. For all considered variables, the standardized mean difference between the two groups is well below the threshold of 0.20 recommended by Imbens and Rubin (2015). Importantly, treated and control units are also balanced in terms of pre-treatment political outcome variables which were not part of the matching process. In both groups, approximately 7% of individuals were members of a political party in 2008. Nearly 2% newly registered as a member in the election cycle between 2005 and 2008, and 0.1% ran for local councilor in the 2008 election. [^20] Appendix Table A.3 also shows that there are no differences in the ideology, local incumbency, and union affiliation of the parties that treated and control observations affiliated with or ran for before the layoff. The validity of our DiD design does not require that treated and control individuals are similar in levels, but covariate balance makes it more plausible that both groups exhibit similar trends in counterfactual outcomes (parallel-trends assumption).
@@
  • Note: The table reports the average characteristics of treated workers who are displaced from a mass layoff firm (column (1)) and matched control workers who are not displaced in the same calendar year (column (2)), and the standardized difference between the two groups (column (3)). The sample only includes workers with a unique name within the country. Appendix Table A.3 reports results for additional pre-layoff political outcomes split by party characteristics.
4.2 Estimation model
We estimate the following dynamic DiD (event-study) model:
@@ 4.2 Estimation model
where $i$ denotes workers and $c$ denotes four-year municipal election cycles (2000, 2004, 2008, 2012, 2016, 2020). Our main outcomes, $Y_{ic}$, are binary variables capturing new party memberships and candidacies for local councils in a given election cycle. [^21] The membership dummy captures new party registrations within the four-year electoral cycles, i.e., between October 1996 and October 2000, November 2000 and October 2004, etc. Although we have information on the exact date individuals affiliate with a political party, we focus on cycle-level information since new affiliations are highly concentrated around election years (see Appendix Figure A.1). Later on, we present a year-level analysis, which allows us to test for pre-trends closer to the layoff event and to assess political responses in the shorter run. The effects of interest are captured by $\delta_{c}$, the coefficients for interactions between a treatment group dummy, $Treat_{i}$, and dummies for each election cycle, $Cycle_{c}$. For $c=\{2012,2016,2020\}$, the parameters identify the dynamic treatment effects (relative to the baseline period $c=2008$), and for $c=\{2000,2004\}$ they check whether outcomes evolved in parallel in the pre-treatment period. [^22] Note that our estimation model relies on a single treatment period, as all treated workers were laid off in the 2012 cycle. Thus, the model is not susceptible to weighting issues that arise in staggered difference-in-differences models with heterogeneous treatment effects (de Chaisemartin and D’Haultfœuille, 2020; Goodman-Bacon, 2021; Sun and Abraham, 2021). We also estimate a static version of the DiD model:
$Y_{ic}=\alpha+\beta Treat_{i}+\delta_{DiD}(Treat_{i}\times\mathbbm{1}[c\geq 2012])+\mu_{m(i)k(i)c}+\epsilon_{ic}.$ (2)
which summarizes the average treatment effects over all periods in the coefficient $\delta_{DiD}$. In both models, we include municipality $\times$ two-digit industry $\times$ cycle fixed effects, $\mu_{m(i)k(i)c}$, which flexibly control for local and industry shocks that may simultaneously cause mass layoffs and changes in political behavior. Importantly, the municipality $\times$ cycle fixed effects also hold constant the local political environment and allow us to separate individual-level supply-side effects from potential effects of mass layoffs on the demand for politicians among the local electorate.
5 Effects of Job Loss on Political Entry

### v3

@@ 2.2 Local councilors
Brazil has a highly decentralized political system, with municipal governments managing substantial budgets to provide public services, such as local infrastructure, transportation, education, and health care. Each of Brazil’s 5,568 municipalities is governed by a mayor (prefeito) and a council of local legislators (vereadores). Every four years, during the municipal elections, citizens cast one vote for a mayoral candidate and one vote for a council candidate. [^6] Local elections are held at the same time for all municipalities across the country, usually in October. Elections for the President, governors, and federal and state legislators also take place every four years but are staggered by two years relative to local elections. Local councilors are elected based on a system of open-list proportional representation and serve four-year terms, with no term limits. The size of local councils varies depending on population, with a minimum of 9 councilors elected in municipalities with up to 15,000 inhabitants and a maximum of 55 councilors elected in municipalities with more than 8 million inhabitants. Local legislators are responsible for drafting laws, voting bills, proposing public works, and monitoring the actions of the executive. In addition, they must approve the municipal budgets and thereby have great influence over the delivery of public goods.
As in many other contexts, serving as a local councilor in Brazil can be financially rewarding. [^7] According to PoliticalSalaries.com, a not-for-profit open data project, being a legislator is financially attractive in many countries, as measured by legislator salaries as a percentage of GDP per capita. In Brazil, legislators earn salaries equivalent to 444% of GDP per capita, a ratio similar to other upper-middle-income countries, such as Colombia (490%) and South Africa (443%). The wages of councilors are substantial on average – amounting to about 2.5 times the average wage in their municipalities. Councilor wages also vary strongly by municipality size – being capped at 20% of state deputies’ wages in municipalities with less than 10,000 inhabitants and at 75% in municipalities with more than 500,000 inhabitants. Importantly, serving as a councilor is not a full-time activity and allows legislators to retain their private-sector jobs after being elected. On average, legislators need to be present in the council only four days per month (Ferraz and Finan, 2011). Besides the salaries received by elected councilors, patronage channels can also generate returns for candidates who are not elected. Colonnelli et al. (2020) find that non-elected councilor candidates are about 15 percentage points more likely to be employed in the public sector and receive 40% higher earnings in the formal economy if their supported party comes into power.
@@
How large are the financial costs associated with candidacies for local council? To finance their campaigns, candidates rely on party funds, financial contributions from individuals and firms, and their own resources. [^10] Campaign spending limits were introduced in 2015 (Avis et al., 2022). In addition to banning financial contributions from private firms, the electoral law also set spending limits for political campaigns based on the population of the municipality. Since 2019, candidates for local council are only allowed to fund up to 10% of their municipality’s spending limit with their own resources, ensuring more equitable running conditions. We collected data from the Superior Electoral Court (Tribunal Superior Eleitoral, TSE) on all campaign contributions received by councilor candidates in the electoral cycles 2004 to 2020. Appendix Table A.1 presents detailed summary statistics for each year. Across all years, the median candidate raised a total of 1,587 BRL (392 USD), with only 168 BRL (34 USD) coming from own resources. Only about half of all candidates use any private funds for the campaign. These figures suggest that personal financial contributions play a limited role in local campaigns, suggesting that the financial barriers to running for councilor are relatively low.
@@ 2.3 Theoretical considerations
We now introduce a simple economic framework to illustrate how a layoff may affect an individual’s likelihood of becoming a party member or councilor candidate. [^11] For economic models of candidate selection, see, for example, Black (1972), Caselli and Morelli (2004), and Messner and Polborn (2004). Consider an individual who decides how to divide her non-leisure time between pursuing a career in the private sector or in politics. Let $h^{P}$ denote the share of time devoted to politics. If the individual does not engage in politics ($h^{P}=0$), she earns a market hourly wage of $w^{M}$. When entering politics ($0<h^{P}\leq 1$), she foregoes a share of her earnings that is proportional to $h^{P}$. [^12] Since neither party membership nor local councilor is a full-time political commitment, we consider the possibility of simultaneously working in the private sector, albeit with a reduced number of hours. At the same time, political engagement provides her with an array of expected benefits, $b^{P}$, which increase with the time devoted to political activities. These benefits include monetary rewards, such as councilor salaries when elected to office (Ferraz and Finan, 2011), public sector employment opportunities resulting from political connections to the winning party or candidate (Brollo et al., 2017; Colonnelli et al., 2020; Barbosa and Ferreira, 2023), and improved employment prospects in the private labor market under co-partisan employers (Colonnelli et al., 2022). Non-monetary rewards, such as ego rents or prosocial benefits, may also play a role (Gulzar and Khan, 2024). Finally, to engage in politics, the individual incurs entry costs $c^{P}$ (e.g., campaign costs or membership registration fees).
@@
An individual who is currently employed in the private sector decides to enter politics if the benefits of engaging in politics are greater than the benefits of focusing exclusively on market work, i.e., if $w^{M}(1-h^{P})+b^{P}h^{P}-c^{P}\geq w^{M}$, or equivalently, if the net benefits are positive:
$\pi_{\text{pre}}=h^{P}(b^{P}-w^{M})-c^{P}\geq 0.$
If the individual is laid off, she experiences an earnings loss that may result from both lower employment prospects and a wage penalty upon reemployment. Let $\lambda$ denote the reemployment probability and $\tilde{w}^{M}$ the new potential market wage, such that the expected labor income is given by $\lambda\tilde{w}^{M}$. If the individual finds reemployment, her expected benefits from engaging in politics become $\tilde{w}^{M}(1-h^{P})+b^{P}h^{P}-c^{P}$. If not reemployed, she earns no labor income and can devote her entire time to politics, yielding benefits of $b^{P}-c^{P}$. [^13] In practice, unemployed individuals may not allocate their entire non-leisure time to politics, but also engage in other activities such as job search. The model’s qualitative predictions remain unchanged as long as the time devoted to politics while unemployed exceeds the level of political time investment under reemployment – that is, it exceeds $h^{P}$. The individual will enter politics if its returns are greater than the expected value of pure market work, i.e., if $\lambda\left(\tilde{w}^{M}(1-h^{P})+b^{P}h^{P}-c^{P}\right)+(1-\lambda)\left(b^{P}-c^{P}\right)\geq\lambda\tilde{w}^{M}$, which leads to the following condition on net benefits:
$\pi_{\text{post}}=\lambda\left(h^{P}(b^{P}-\tilde{w}^{M})-c^{P}\right)+(1-\lambda)\left(b^{P}-c^{P}\right)\geq 0.$
@@
To assess the effect of a layoff on the likelihood of entering politics, we can compare the net benefits before and after the layoff:
$\pi_{\text{post}}-\pi_{\text{pre}}=h^{P}(w^{M}-\lambda\tilde{w}^{M})+(1-h^{P})(1-\lambda)b^{P}.$
With $w^{M}\geq\tilde{w}^{M}\geq 0$, $b^{P}\geq 0$, $0\leq h^{P}\leq 1$, and $0\leq\lambda\leq 1$, this difference is weakly positive, implying that layoffs may induce more individuals to enter politics. The change in net benefits resulting from the layoff has two components. First, it is proportional to the expected income loss, $w^{M}-\lambda\tilde{w}^{M}$, weighted by the share of time not allocated to private-sector work, $h^{P}$. This term captures the decline in monetary opportunity costs arising from the layoff: as earnings potential in the private sector falls, individuals have to give up less income when engaging in political activities, particularly those that are time-intensive. To empirically test the relevance of this mechanism, we will analyze heterogeneity in treatment effects by individuals’ expected income losses due to the layoff, which are predicted based on a rich set of individual and local labor market characteristics.
@@
Second, the effects increase with the expected benefits of political engagement during unemployment, $(1-\lambda)b^{P}$, weighted by the time not already allocated to politics, $1-h^{P}$. Intuitively, this term reflects the notion that, when remaining unemployed, individuals can reallocate time that would otherwise be spent in market work towards political activities, thereby gaining additional political benefits. To shed light on one of the monetary returns from engaging with politics, we will investigate whether the treatment effects are larger in localities where councilors earn higher salaries. [^14] In addition to monetary returns, political involvement may also provide policy payoffs. In citizen-candidate models (Osborne and Slivinski, 1996; Besley and Coate, 1997), political candidates represent particular policies and care about winning per se. As argued by Hall (2019) and Thomsen (2017), the policy returns may be especially relevant for ideologically extreme candidates. Thus, we will examine potential heterogeneity in the effects on memberships and candidacies in parties of different ideological positions. Moreover, we will test the role of time resources by exploiting discontinuities in eligibility for UI benefits that increase laid-off workers’ incentives to remain unemployed.
Note that the relative importance of both channels – reduced income opportunities and increased time availability – is determined by the time intensity of political engagement, $h^{P}$. For forms of engagement that require substantial time investments, such as running for office, the income-loss channel is likely more important. By contrast, for lower-commitment forms of participation, such as party membership, the time-availability channel may play a larger role.
@@
The framework also provides predictions about the potential implications of layoffs for the quality of individuals who select into politics. Given that the participation effects of layoffs are predicted to be larger for individuals with higher income losses and higher expected benefits from political engagement, the consequences for selection depend on how these factors are related to individual competence. High-ability individuals – such as those with higher education – tend to earn higher wages in the private sector. Thus, a layoff may represent a larger shock to their opportunity costs, leading to a more pronounced increase in their incentives to enter politics. In addition, the expected benefits of political engagement may vary with individuals’ competence. For instance, the returns to running for office rise with the probability of winning, which may be higher for more capable individuals. Moreover, the expected likelihood of gaining access to a public-sector job through political connections may depend on ability, either negatively if political support acts as a substitute for individual competence (Colonnelli et al., 2020) or positively if politicians use their discretionary power to appoint more competent individuals (Brollo et al., 2017; Bazzi and Labanca, 2023).
@@ 3 Data
In our analysis, we combine detailed information on the universe of party members and local political candidates with administrative data on the population of formal employees in Brazil. This section describes the data sources as well as the steps we follow to construct and merge our estimation sample. By having identified records on individuals’ employment histories as well as political outcomes, we can uniquely identify whether individuals’ political behavior changes after they lose their jobs. Additionally, detailed information on individual, municipality, and party characteristics allows us to disentangle potential mechanisms.
3.1 Political outcomes
@@
3.2 Labor market data
We obtain matched employer-employee data covering the universe of formal employees between 2004 and 2018 from the Relação Anual de Informações Sociais (RAIS), an administrative database collected annually by the Ministry of Labor and Employment. [^15] In 2015, formally employed individuals accounted for approximately 55% of the Brazilian labor force (Derenoncourt et al., 2021). The database was created in 1975 for statistical and administrative purposes and contains detailed information such as hiring and separation date, average monthly earnings, contracted hours, occupation, firm’s location and industry, and worker’s socio-demographic characteristics. For our analysis, we make use of the identified dataset, containing individuals’ names, taxpayer registry numbers (CPF), as well as firms’ tax identification numbers (Cadastro Nacional da Pessoa Jurídica, CNPJ). This enables us to identify individuals’ employment histories over time and across firms.
@@ 3.3 Merging political and labor market data
Appendix Table A.2 compares the characteristics of laid-off individuals in our estimation sample who have a unique and non-unique name, respectively. Those with a unique name are slightly positively selected (9% higher earnings, 8% more years of education, 0.9 percentage points more likely to be a manager), and are more likely to be female (8.1 percentage points) and white (5.3 percentage points). Given that we merge our second political outcome – local candidacies – using the CPF, we have this information for individuals with and without a unique name. Before the layoff, we do not observe differences in the likelihood of running for councilor across both groups. While we focus on the sample of individuals with unique names (in which we have merged political members and candidates) in our baseline specification, we also analyze the external validity of the results by estimating layoff effects on candidacies in the sample of individuals with non-unique names.
4 Effects of Job Loss on Political Entry
4.1 Empirical strategy
@@
The main difficulty in estimating the effect of layoffs on individuals’ entry into politics is the fact that layoffs are not random. On the one hand, individual shocks could simultaneously influence both political participation and the likelihood of losing a job, leading to a spurious relationship. On the other hand, individuals’ political behavior could also affect their probability of being dismissed. Colonnelli et al. (2022) show, for instance, that political discrimination affects labor market outcomes in Brazil. To address these identification challenges, we exploit the timing of firm-level mass layoffs as a source of plausibly exogenous variation in individual-level job loss using a difference-in-difference design in a matched sample of workers. The timing of mass layoffs is arguably orthogonal to individual workers’ political behavior and other confounding worker-level shocks, and has been widely used to estimate the effects of job loss on various outcomes. In particular, our empirical strategy closely follows recent papers that study the effect of dismissals in Brazil (Bhalotra et al., 2025; Britto et al., 2022b, a; Amorim et al., 2025). This subsection describes how we construct the estimation sample and then presents the estimation model.
@@
Sample selection. We restrict our sample to workers aged 25 to 50 with a full-time (i.e., at least 30 hours per week), open-ended, private-sector job of at least six months tenure. Among those, the treatment group comprises all workers displaced without just cause from a mass layoff firm between 2009 and 2011.171717We exclude individuals who are recalled by their initial employer within five years after the layoff. In cases where individuals experienced multiple mass layoffs between 2009 and 2011, we keep only the first one. This treatment period allows us to study long-run effects over three election cycles (12 years) after the layoff, and also to check for pre-trends over two election cycles (8 years) before the layoff. In our baseline specification, we define a mass layoff firm as a firm with at least 30 workers dismissing at least 30% of its workforce in a given year, and later we show that our results are robust to various thresholds of firm size and layoff share. [^18] We always drop firms from the mass layoff sample if they reallocate under a new identifier, specifically if at least 30% of their dismissed workers transition to the same new firm identifier in the following year.
@@
  • Note: The table reports the average characteristics of treated workers who are displaced from a mass layoff firm (column (1)) and matched control workers who are not displaced in the same calendar year (column (2)), and the standardized difference between the two groups (column (3)). The sample only includes workers with a unique name within the country. Appendix Table B.1 reports results for additional pre-layoff political outcomes split by party characteristics.
Table 1 presents summary statistics for treated and control individuals and shows that both groups are similar in socio-demographic, job, and firm characteristics. For all considered variables, the standardized mean difference between the two groups is well below the threshold of 0.20 recommended by Imbens and Rubin (2015). Importantly, treated and control units are also balanced in terms of pre-treatment political outcome variables which were not part of the matching process. In both groups, approximately 7% of individuals were members of a political party in 2008. Nearly 2% newly registered as a member in the election cycle between 2005 and 2008, and 0.1% ran for local councilor in the 2008 election. [^21] Appendix Table B.1 also shows that there are no differences in the ideology, local incumbency, and union affiliation of the parties that treated and control observations affiliated with or ran for before the layoff. The validity of our DiD design does not require that treated and control individuals are similar in levels, but covariate balance makes it more plausible that both groups exhibit similar trends in counterfactual outcomes.
Estimation model. As our main specification for the effect of job loss on political participation, we estimate the following dynamic DiD (event-study) model:
@@
where $i$ denotes workers and $c$ denotes four-year municipal election cycles (2000, 2004, 2008, 2012, 2016, 2020). Our main outcomes, $Y_{ic}$, are binary variables capturing new party memberships and candidacies for local councils in a given election cycle. The effects of interest are captured by $\delta_{c}$, the coefficients for interactions between a treatment group dummy, $Treat_{i}$, and dummies for each election cycle, $Cycle_{c}$. For $c=\{2012,2016,2020\}$, the parameters identify the dynamic treatment effects (relative to the baseline period $c=2008$), and for $c=\{2000,2004\}$ they check whether outcomes evolved in parallel in the pre-treatment period. We also estimate a static version of the DiD model:
$Y_{ic}=\alpha+\beta Treat_{i}+\delta_{DiD}(Treat_{i}\times\mathbbm{1}[c\geq 2012])+\mu_{m(i)k(i)c}+\epsilon_{ic},$ (2)
which summarizes the average treatment effects over all periods in the coefficient $\delta_{DiD}$. In both models, we include municipality $\times$ two-digit industry $\times$ cycle fixed effects, $\mu_{m(i)k(i)c}$, which flexibly control for local and industry shocks that may simultaneously cause mass layoffs and changes in political behavior. Importantly, the municipality $\times$ cycle fixed effects also hold constant the local political environment and allow us to separate individual-level supply-side effects from potential effects of mass layoffs on the demand for politicians among the local electorate. [^22] Note that the estimation models (1) and (2) rely on a single treatment period, as all treated workers were laid off in the 2012 cycle. Thus, our main specification is not susceptible to weighting issues that arise in staggered difference-in-differences models with heterogeneous treatment effects (de Chaisemartin and D’Haultfœuille, 2020; Goodman-Bacon, 2021; Sun and Abraham, 2021).
While our main specification analyzes party memberships and candidacies in each election cycle, for party memberships and labor market outcomes we can also exploit yearly variation. This allows us to test for pre-trends closer to the layoff event and assess responses in the shorter run. [^23] We have information on the exact date individuals affiliate with a political party. Still, we focus on cycle-level variation in our main specification, since new affiliations are highly concentrated around election years (see Appendix Figure A.2), and use the year-level model as a supplementary analysis of effect dynamics. For labor market outcomes, we always show results with the year-level model, because we do not observe outcomes for all election cycles between 2000 and 2020 (we only have access to RAIS between 2004 and 2018). We estimate the following event-study model for outcomes measured in each year relative to the layoff:
where $i$, $t$, and $t^{*}$ refer to workers, calendar years, and layoff years, respectively. [^24] For control workers, we assign the layoff year of their matched treated worker. $\delta_{k}$ capture the dynamic treatment effects for each event year after the layoff $k$. The model includes calendar year $\times$ layoff year fixed effects, $\lambda_{tt^{*}}$, to ensure that the identifying variation only comes from comparing treated to control individuals who have been matched in the same year. [^25] This stacking approach avoids the “forbidden comparisons” between late and early-treated individuals that can cause negative weights under heterogeneous treatment effects (Cengiz et al., 2019; Dube et al., 2023).